40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13051     Entry time: Wed Jun 7 17:45:11 2017     In reply to: 13048     Reply to this: 13052
Author: gautam 
Type: Update 
Category: ASS 
Subject: Y-arm coil driver electronics investigation 

I repeated the test of driving C1:SUS-<Optic>_<coil>_EXC individually and measuring the transfer function to C1:SUS-<Optic>_OPLEV_PERROR for Optic in (ITMX, ITMY, ETMX, ETMY, BS), coil in (LLCOIL, LRCOIL, ULCOIL, URCOIL). 

There seems to be a few dB imbalance in the coils in both ETMs, as well as ITMX. ITMY and the BS seem to have pretty much identical TFs for all the coils - I will cross-check using OPLEV_YERROR, but is there any reason why we shouldn't adjust the gains in the coil output (not output matrix) filter banks to correct for this observed imbalance? The Oplev calibrations for the various optics are unknown, so it may not be fair to compare the TFs between optics (I guess the same applies to comparing TF magnitudes from coil to OPLEV_PERROR and OPLEV_YERROR, perhaps we should fix the OL calibrations before fiddling with coil gains...)

The anomalous behaviour of ITMY_UL (10dB greater than the others) was traced down to a rogue x3 gain in the filter module indecision. This has been removed, and now Y arm ASS works fine (with the original dither servo settings). X arm dither still doesn't converge - I double checked the digital filters and all seems in order, will investigate the analog part of the drive electronics now.

 

Attachment 1: CoilTFs.pdf  163 kB  | Hide | Hide all
CoilTFs.pdf CoilTFs.pdf CoilTFs.pdf
ELOG V3.1.3-